Aspirin has the formula C9H804. How many atoms Of oxygen are there in a tablet weighing
360 mg?
1
Expert's answer
2020-11-18T13:40:05-0500

C9H8O4 = 4 O atoms

180g of aspirin = 4 mole of O atom

1 g of aspirin = 4 /180 mole of O atom

360mg of aspirin = 4/180 × 0.36 mole of O atom

= 0.008 moles of O atom

1 mole of O atom = 6.02 × 1023 atoms

0.008 mole of O = 0.008 × 6.02 × 1023 atoms

= 4.82 × 1021 atoms